1. An appliance retailer purchased a small vacuum cleaner for $59.99. He plans to take a 33% markup. What will the selling price be? (Value 5 pts)

2. If a retailer sells a coat for $139.99. What was his markup if he bought the coat for $80.00. (Value 10 pts)

Markup in actual dollars ________

Express markup as a % of cost _________

Express markup as a % of selling price ________
Answer #1

1. Selling price calculation

Cost of purchase of vacuum cleaner = 59.99

Markup =33

Selling price = Cost of purchase + (cost of purchase × markup)

Selling price= 59.99 + (59.99×0.33)= 79.79

Selling price = Cost of purchase × (1+markup) formula is another formula

Selling price = 59.99× 1.33= 79.79

2. Markup in actual dollars= selling price - cost price

Markup in actual dollars = 139.99 - 80= 59.99

Express markup as a % of cost= markup in dollars / cost price

Express markup as a % of cost = 59.99/80= 74.99

Express markup as a % of selling price= markup in dollars / selling price

Express markup as a % of selling price = 59.99/139.99= 42.85%